# A Host–Kra $\mathbb{F}_{2}^{\omega}$-system of order $5$ that is not Abramov of order $5$, and non-measurability of the inverse theorem for the $U^{6}(\mathbb{F}_{2}^{n})$ norm

### v2

@@ A.3. Polynomials
Let $P\colon G\to H$ be a map from a filtered abelian group $G$ to an abelian group $H$. If $k\geq-1$, then $P$ is a polynomial of degree at most $k$ if and only if $d^{k+1}P=0$. In particular, for $k\geq 0$, we see that $P$ is a polynomial of degree at most $k$ if and only if $dP$ is a polynomial of degree at most $k-1$.
As one application of this proposition, we have the following familiar-looking result about multiplication of polynomials (cf. [24, Exercise 1.6.10]):
@@
for all $x_{1},\dots,x_{n}\in\mathbb{F}_{2}$ and some $10\leq\alpha<$ 1 and some integers $0\leq c_{i_{1},\dots,i_{k}}<2^{d+1-k}$, where $|x|\coloneqq 1_{x=1}$. The coefficients $\alpha$ and $c_{i_{1},\dots,i_{k}}$ are uniquely determined. Indeed we have
@@
This follows from [26, Lemma 1.7(iii)], with the latter identities following from a routine calculation. There is an analogous classification of polynomials in other characteristics than $2$, but we will only need the characteristic two theory here. ∎
@@
for all $d\geq 0$, refining Proposition A.22 in this case; thus, every polynomial $P$ of degree $d$ can be expressed in the form $P=2Q$ for some polynomial $Q$ of degree $d+1$, and conversely if $Q$ is of degree $d+1$ then $2Q$ is of degree $d$; see [26, Lemma 1.7(v)]. In a similar spirit, we have
@@
for all $x_{1},\dots,x_{n}\in\mathbb{F}_{2}$ and some $10\leq\alpha<$ 1 and some integers $0\leq c_{i_{1},\dots,i_{k}}<2^{d+1-k}$, where $|x|=1_{x=1}$. We then define $Q(x_{1},\dots,x_{n})$ explicitly by the formula
@@ A.4. $p$-homogeneous nilspaces
[4, Definitions 1.2, 3.1] Let $p$ be a prime. A nilspace $X$ is said to be $p$-homogeneous if, whenever $n\geq 0$ and $f\colon{\mathcal{D}}^{1}(\mathbb{Z}^{n})\to X$ is a nilspace morphism, then the periodization $\tilde{f}\colon{\mathcal{D}}^{1}(\mathbb{F}_{p}^{n})\to X$, defined by restricting $f$ to $\{0,\dots,p-1\}^{n}$ and then extending periodically, is also a nilspace morphism.
@@
In [4, Remark 3.3] it is noted that $p$-homogeneity implies weak $p$-homogeneity, and that the two concepts are equivalent when $p=2$. In [4, Question 3.4] it is posed as an open question whether these two concepts are equivalent for $p>2$; we do not address this question here. From this remark, we see that $X$ is $2$-homogeneous (or equivalently, weakly $2$-homogeneous), if and only if we have a bijection
@@
There is an analogue of this result for general $p$, but it is more difficult to prove; see [4, Proposition 3.12].
@@
since $02e_{i}=$ 0 (cf. (4)); since $Z$ is assumed to be an elementary abelian $2$-group, we thus also have
@@
See [4, Theorem 1.4]. In fact the ergodicity hypothesis can be dropped here, but we will not need to use this fact. ∎
@@ Appendix B Deducing the strong inverse conjecture from the BTZ conjecture
Fix $p,k,\eta,\varepsilon()$; all quantities below are permitted to depend on these parameters. Suppose for contradiction that Conjecture 1.2 was true, but Conjecture 1.3 failed for the indicated choice of $\eta,\varepsilon()$. Without loss of generality we may assume $\varepsilon(m)\leq\frac{1}{m}$ (for instance). Then for every $M$, there exists $G=\mathbb{F}_{p}^{n}$ for some $n=n_{M}$ and a function $f=f_{M}\colon G\to{\mathcal{D}}$ with $\|f\|_{U^{k+1}(G)}\geq\eta$, such that if $h_{1},\dots,h_{M}\in G$ are chosen independently and uniformly at random, then with probability greater than $1/2$, there does not exist $1\leq m\leq M$ and $P\in\mathrm{Poly}^{k}(G)$ and a function $F\colon\mathbb{C}^{\mathbb{F}_{p}^{M}}\to\mathbb{C}$ of Lipschitz constant at most $M$, such that
@@
such that for any $G=\mathbb{F}_{p}^{n}$ and $f\colon G\to{\mathcal{D}}$, if $v_{1},v_{2},v_{3},\dots\in G$ are chosen uniformly and independently at random, then with probability at least $1-\varepsilon_{0}$, the following “accurate sampling” statement holds: for every sequence
@@
and every Lipschitz $F\colon{\mathcal{D}}^{\{0,1\}^{k+1}\times\mathbb{F}_{p}^{r_{0}}}\to\mathbb{C}$, we have
@@
See [1, Proposition 3.13] (with some mild relabeling, for instance replacing $k$ by $k+1$). In that proposition the sampling property was only asserted to hold with positive probability, but an inspection of the proof shows that it can be established with probability at least $1-\varepsilon_{0}$ for any fixed $\varepsilon_{0}>0$. ∎
@@
We fix this data for each $M$. Following [26], we now introduce the universal Furstenberg space $X\coloneqq{\mathcal{D}}^{\mathbb{F}_{p}^{\omega}}$ of functions $\zeta\colon\mathbb{F}_{p}^{\omega}\to{\mathcal{D}}$ with the product $\sigma$-algebra and shift action
@@
Let $f_{\infty}\colon X\to{\mathcal{D}}$ be the coordinate function
@@
[…]
for any $l,m\geq 1$, any $\vec{a}_{j}=(a_{j,i})_{i=1}^{\infty}\in\mathbb{F}_{p}^{\omega}$ for $j=1,\dots,l$, and any continuous function $F\colon{\mathcal{D}}^{\ell}\to\mathbb{C}$. This allows us to pass back and forth between integral expressions on $X$ (using the measure $\mu_{M}$) and combinatorial averages on $\mathbb{F}_{p}^{n_{M}}$.
@@
[…]
(say). Let $c:\mathbb{R}^{+}\to\mathbb{R}^{+}$ be a decreasing function to be chosen later (depending on $k,p$) such that $c(\varepsilon)\to 0$ decays to zero sufficiently rapidly as $\varepsilon\to 0$. Then, as $X$ is generated by $f_{\infty}$ and its shifts, we see that there exists a natural number $M_{0}$ and shifts $\vec{b}_{1},\dots,\vec{b}_{M_{0}}in\mathbb{F}_{p}^{\omega}$, and a Lipschitz function $F\colon{\mathcal{D}}^{M_{0}}\to{\mathcal{D}}$ of Lipschitz constant at most $M_{0}$ such that
[…]
@@
[…]
@@
[…]
@@
[…]

### v3

@@ A.3. Polynomials
Let $P\colon G\to H$ be a map from a filtered abelian group $G$ to an abelian group $H$. If $k\geq-1$, then $P$ is a polynomial of degree at most $k$ if and only if $d^{k+1}P=0$. In particular, for $k\geq 0$, we see that $P$ is a polynomial of degree at most $k$ if and only if $dP$ is a polynomial of degree at most k1k-1.
As one application of this proposition, we have the following familiar-looking result about multiplication of polynomials (cf. [25, Exercise 1.6.10]):
@@
for all $x_{1},\dots,x_{n}\in\mathbb{F}_{2}$ and some $0\leq\alpha<1$ and some integers $0\leq c_{i_{1},\dots,i_{k}}<2^{d+1-k}$, where $|x|\coloneqq 1_{x=1}$. The coefficients $\alpha$ and $c_{i_{1},\dots,i_{k}}$ are uniquely determined. Indeed we have
@@
This follows from [27, Lemma 1.7(iii)], with the latter identities following from a routine calculation. There is an analogous classification of polynomials in other characteristics than $2$, but we will only need the characteristic two theory here. ∎
@@
for all $d\geq 0$, refining Proposition A.22 in this case; thus, every polynomial $P$ of degree $d$ can be expressed in the form $P=2Q$ for some polynomial $Q$ of degree $d+1$, and conversely if $Q$ is of degree $d+1$ then $2Q$ is of degree $d$; see [27, Lemma 1.7(v)]. In a similar spirit, we have
@@
for all $x_{1},\dots,x_{n}\in\mathbb{F}_{2}$ and some $0\leq\alpha<1$ and some integers $0\leq c_{i_{1},\dots,i_{k}}<2^{d+1-k}$, where $|x|=1_{x=1}$. We then define $Q(x_{1},\dots,x_{n})$ explicitly by the formula
@@ A.4. $p$-homogeneous nilspaces
[5, Definitions 1.2, 3.1] Let $p$ be a prime. A nilspace $X$ is said to be $p$-homogeneous if, whenever $n\geq 0$ and $f\colon{\mathcal{D}}^{1}(\mathbb{Z}^{n})\to X$ is a nilspace morphism, then the periodization $\tilde{f}\colon{\mathcal{D}}^{1}(\mathbb{F}_{p}^{n})\to X$, defined by restricting $f$ to $\{0,\dots,p-1\}^{n}$ and then extending periodically, is also a nilspace morphism.
@@
In [5, Remark 3.3] it is noted that $p$-homogeneity implies weak $p$-homogeneity, and that the two concepts are equivalent when $p=2$. In [5, Question 3.4] it is posed as an open question whether these two concepts are equivalent for $p>2$; we do not address this question here. From this remark, we see that $X$ is $2$-homogeneous (or equivalently, weakly $2$-homogeneous), if and only if we have a bijection
@@
There is an analogue of this result for general $p$, but it is more difficult to prove; see [5, Proposition 3.12].
@@
since $2e_{i}=0$ (cf. (4)); since $Z$ is assumed to be an elementary abelian $2$-group, we thus also have
@@
See [5, Theorem 1.4]. In fact the ergodicity hypothesis can be dropped here, but we will not need to use this fact. ∎
@@ Appendix B Deducing the strong inverse conjecture from the BTZ conjecture
Fix $p,k,\eta,\varepsilon()$; all quantities below are permitted to depend on these parameters. Suppose for contradiction that Conjecture 1.2 was true, but Conjecture 1.3 failed for the indicated choice of $\eta,\varepsilon()$. Without loss of generality we may assume $\varepsilon(m)\leq\frac{1}{m}$ (for instance). Then for every $M$, there exists $G=\mathbb{F}_{p}^{n}$ for some $n=n_{M}$ and a function $f=f_{M}\colon G\to\mathbb{D}$ with $\|f\|_{U^{k+1}(G)}\geq\eta$, such that if $h_{1},\dots,h_{M}\in G$ are chosen independently and uniformly at random, then with probability greater than $1/2$, there does not exist $1\leq m\leq M$ and $P\in\mathrm{Poly}^{k}(G)$ and a function $F\colon\mathbb{C}^{\mathbb{F}_{p}^{M}}\to\mathbb{C}$ of Lipschitz constant at most $M$, such that
@@
such that for any $G=\mathbb{F}_{p}^{n}$ and $f\colon G\to\mathbb{D}$, if $v_{1},v_{2},v_{3},\dots\in G$ are chosen uniformly and independently at random, then with probability at least $1-\varepsilon_{0}$, the following “accurate sampling” statement holds: for every sequence
@@
and every Lipschitz $F\colon\mathbb{D}^{\{0,1\}^{k+1}\times\mathbb{F}_{p}^{r_{0}}}\to\mathbb{C}$, we have
@@
See [26, Proposition 3.13] (with some mild relabeling, for instance replacing $k$ by $k+1$). In that proposition the sampling property was only asserted to hold with positive probability, but an inspection of the proof shows that it can be established with probability at least $1-\varepsilon_{0}$ for any fixed $\varepsilon_{0}>0$. ∎
@@
We fix this data for each $M$. Following [26], we now introduce the universal Furstenberg space $X\coloneqq\mathbb{D}^{\mathbb{F}_{p}^{\omega}}$ of functions $\zeta\colon\mathbb{F}_{p}^{\omega}\to\mathbb{D}$ with the product $\sigma$-algebra and shift action
@@
Let $f_{\infty}\colon X\to\mathbb{D}$ be the coordinate function
@@
(67) $\begin{split}&\int_{X}F(T_{\vec{a}_{1}}f_{\infty},\dots,T_{\vec{a}_{l}}f_{\infty})\ d\mu_{M}\\ &\quad=\mathbb{E}_{x\in\mathbb{F}_{p}^{n_{M}}}F\left(f_{M}\left(x+\sum_{i=1}^{\infty}a_{1,i}v_{i,M}\right),\dots,f_{M}\left(x+\sum_{i=1}^{\infty}a_{l,i}v_{i,M}\right)\right)\end{split}$
for any $l,m\geq 1$, any $\vec{a}_{j}=(a_{j,i})_{i=1}^{\infty}\in\mathbb{F}_{p}^{\omega}$ for $j=1,\dots,l$, and any continuous function $F\colon\mathbb{D}^{\ell}\to\mathbb{C}$. This allows us to pass back and forth between integral expressions on $X$ (using the measure $\mu_{M}$) and combinatorial averages on $\mathbb{F}_{p}^{n_{M}}$.
@@
$\left|\int_{X}f_{\infty}e(-P)\ d\mu\right|>\frac{3}{m}$
(say). Let $c:\mathbb{R}^{+}\to\mathbb{R}^{+}$ be a decreasing function to be chosen later (depending on $k,p$) such that $c(\varepsilon)\to 0$ decays to zero sufficiently rapidly as $\varepsilon\to 0$. Then, as $X$ is generated by $f_{\infty}$ and its shifts, we see that there exists a natural number $M_{0}$ and shifts $\vec{b}_{1},\dots,\vec{b}_{M_{0}}in\mathbb{F}_{p}^{\omega}$, and a Lipschitz function $F\colon\mathbb{D}^{M_{0}}\to\mathbb{D}$ of Lipschitz constant at most $M_{0}$ such that
(68) $\int_{X}|e(P)-F(T_{\vec{b}_{1}}f_{\infty},\dots,T_{\vec{b}_{M_{0}}}f_{\infty})|\ d\mu<c(\varepsilon(m))$
@@
$\left|\int_{X}f_{\infty}\overline{F}(T_{\vec{b}_{1}}f_{\infty},\dots,T_{\vec{b}_{M_{0}}}f_{\infty})\ d\mu\right|>\frac{2}{m}.$
@@
$\left|\int_{X}f_{\infty}\overline{F}(T_{\vec{b}_{1}}f_{\infty},\dots,T_{\vec{b}_{M_{0}}}f_{\infty})\ d\mu_{M}\right|\geq\frac{2}{m}.$
@@
$\mathbb{E}_{\vec{a}_{1}\in\mathbb{F}_{p}^{H_{r_{1}}},\dots,\vec{a}_{k+1}\in\mathbb{F}_{p}^{H_{r_{k+1}}}}\int_{X}|\Delta_{\vec{a}_{1}}\dots\Delta_{\vec{a}_{k+1}}F(T_{\vec{b}_{1}}f_{\infty},\dots,T_{\vec{b}_{M_{0}}}f_{\infty})-1|\ d\mu_{M}\ll c(\varepsilon(m))$
@@
Declarations
Or Shalom was supported by NSF grant DMS-1926686 and ISF grant 3056/21. Terence Tao was supported by a Simons Investigator grant, the James and Carol Collins Chair, the Mathematical Analysis & Application Research Fund, and by NSF grants DMS-1764034 and DMS-2347850. Asgar Jamneshan was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - 547294463.
The authors have no relevant financial or non-financial interests to disclose. On behalf of all authors, the corresponding author states that there is no conflict of interest.